# BayesFLo: Bayesian fault localization of complex software systems

### v2

@@ 2.2 State-of-the-Art and Its Limitations
Existing fault localization methods, as described earlier, are largely deterministic in nature. This includes the work of Niu et al., (2013), who used a tuple relationship tree to capture relationships among all factor combinations based on testing results. For a given test case, this tree lists all considered combinations (tuples) along the branches of the tree, which can then be used to classify which class (faulty or healthy) each tuple belongs to. Nie and Leung, 2011a introduced the idea of a minimal failure-causing schema, defined as the smallest-order factor combination such that all test cases containing them trigger a failure. This schema is then applied for guiding fault localization and subsequent testing. Ghandehari, (2016) and Ghandehari et al., (2018) developed a two-stage combinatorial-testing-based fault localization approach. The key idea is to identify potential failure-inducing factor combinations from test results by eliminating combinations that appear in passed test cases, then rank such combinations based on two proposed “combination suspiciousness” and “environment suspiciousness” metrics. Lekivetz and Morgan, (2018) proposed a deterministic ranking procedure that incorporates a domain-knowledge-guided weighting scheme. A key limitation of such methods is that they are not probabilistic in nature, and thus do not provide the desired probabilistic measure of confidence for how likely a particular combination is a root cause. The above methods also unfortunately do not have publicly available code; we instead make use of the JMP Covering Array Analysis (JMP Statistical Discovery LLC,, 2023) as the “state-of-the-art” approach, which has integrated such methods either in full or in part.
We first provide a brief overview of the fault localization analysis from the JMP Covering Array module. Similar to Ghandehari et al., (2018), this analysis first removes all combinations that have been cleared in passed cases; we call these “tested-and-passed” combinations later. The remaining combinations are then ranked in terms of its failure count, i.e., the number of failed test cases for which this combination is present. For combinations with the same failure count, those with the smallest-order combination are treated as more suspicious, following the principle of combination hierarchy in Lekivetz and Morgan, (2018). Here, the ranking of potential root cause combinations via its failure count is quite intuitive, as combinations that show up in more failed test cases should naturally be treated as more suspicious.
For our TCAS case study, Figure 2 shows the top ten combinations (all tied in terms of suspiciousness) returned by the JMP Covering Array Analysis. This analysis highlights several limitations with the state-of-the-art. First, the tied nature of these top ten combinations makes its subsequent investigation very costly. For such complex systems, the debugging of each combination can require thousands of dollars (Bush,, 1990); the budget required for debugging ten combinations can thus be prohibitively expensive. Second, this analysis is deterministic and does not shed light on the probability that a suspicious combination is a root cause. Such a probabilistic quantification of risk is needed for trustworthy fault diagnosis (Zhou et al.,, 2023). A further benefit of a probabilistic analysis is its potential for disentangling the tied combinations noted above. This potential for disentangling effects was noted in Chang, (2019), who observed that a Bayesian framework can help de-alias effects in fractional factorial designs; see also related work by Mak and Wu, (2019). We show later that a carefully-specified Bayesian model can similarly help disentangle the many tied combinations from Figure 2, facilitating timely and cost-efficient fault diagnosis.
Figure 2: Top suspicious combinations from the JMP Covering Array analysis for our motivating TCAS case study, with its corresponding failure counts from test runs.
@@ 3 The BayesFLo Model
We propose a new Bayesian Fault Localization (BayesFLo) framework, which provides a principled probabilistic approach for fault localization via conditioning on test set outcomes. We first present the employed Bayesian model, then show how it permits the integration of domain and structural knowledge within its prior specification. With such knowledge integrated, BayesFLo can then facilitate timely and confident fault localization with limited costly test runs.
Here, it is worth mentioning that, while there may be different “types” of failures (e.g., incorrect or missing outputs), the system has typically undergone a reasonable level of preliminary testing (e.g., unit testing, which eliminates root causes in single units) at time of combinatorial testing. Because of this, there remains only a very limited number of faults for diagnosis (Lekivetz and Morgan,, 2021). As such, we presume in the following only a single failure type at time of BayesFLo analysis. We will briefly discuss an approach for the rare scenario of multiple failure types later in the Conclusion.
@@ 3.1 Prior Specification
We first introduce some notation. Consider a software system (or more broadly, a complex engineering system) with $I\geq 1$ input factors, where a factor $i$ can take on $J_{i}\geq 2$ different levels. A $K$-input combination (with $K\leq I$) is denoted as $(\mathbf{i},\mathbf{j})_{K}$, where $\mathbf{i}=(i_{1},\cdots,i_{K}),i_{1}<\cdots<i_{K}$ is an ordered $K$-vector containing all inputs for this combination, and $\mathbf{j}=(j_{1},\cdots,j_{K}),j_{k}\in\{1,\cdots,J_{i_{k}}\}$ is a $K$-vector indicating the levels of each corresponding input. For example, the 2-input combination of the first factor at level 1 and the second factor at level 2 can be denoted as $(\mathbf{i},\mathbf{j})_{2}$, where $\mathbf{i}=(1,2)$ and $\mathbf{j}=(1,2)$. In the case of $K=1$, i.e., a single input $i$ at level $j$, this may be simplified to $(i,j)$.
Now let $\mathcal{C}_{K}$ denote the set of $K$-input combinations $(\mathbf{i},\mathbf{j})_{K}$ as described above, and let $\mathcal{C}=\cup_{K=1}^{I}\mathcal{C}_{K}$ denote the set of combinations over all orders $K=1,\cdots,I$. Further let $Z_{(\mathbf{i},\mathbf{j})_{K}}\in\{0,1\}$ be an indicator variable for whether the combination $(\mathbf{i},\mathbf{j})_{K}$ is truly a root cause. As this is unknown prior to running test cases, we model each $Z_{(\mathbf{i},\mathbf{j})_{K}}$ a priori as an independent Bernoulli random variable of the form:
@@
It is worth noting the sheer number of input combinations in $\mathcal{C}$ that needs to be considered as potential root causes. Assuming each factor has an equal number of levels $J=J_{1}=\cdots=J_{I}$, one can show that $\mathcal{C}_{K}$ contains $\binom{I}{K}J^{K}$ distinct combinations of order $K$, thus the total number of consider input combinations is $|\mathcal{C}|=$ $\sum_{K=1}^{I}\binom{I}{K}J^{K}$. Even with a moderate number of inputs, say $I=10$, with each having $J=2$ levels, this amounts to $|\mathcal{C}|=59,048$ combinations. As we shall see later in Section 4, the size of $\mathcal{C}$ forms the key bottleneck for Bayesian inference, as the computation of posterior probabilities can require $\mathcal{O}(2^{|C|})$ work; this can thus be infeasible even for small software systems.
@@
In words, the combination root cause probability $p_{(\mathbf{i},\mathbf{j})_{K}}$ is modeled as the product of the root cause probabilities for its component inputs. The key advantage of this form is that it captures structural knowledge on root causes, by embedding the principles of combination hierarchy and heredity (Lekivetz and Morgan,, 2021). These principles capture the structured nature of typical software root causes, and can be seen as extensions of the well-known principles of effect hierarchy and heredity (Wu and Hamada,, 2009), which are widely used for analysis of factorial experiments. The first principle, combination hierarchy, asserts that combinations involving fewer inputs are more likely to be failure-inducing than those involving more inputs. Empirical evidence suggests this principle holds for software systems across various domains (Kuhn et al.,, 2004). To see how our prior (2) captures combination hierarchy, note that by its product form construction, the combination probability $p_{(\mathbf{i},\mathbf{j})_{K}}$ is always less than the probability of any component input $p_{(i_{k},j_{k})}$. Thus, this prior assigns increasingly smaller root cause probabilities on combinations with a higher combination order $K$, thus capturing the desired hierarchy structure. The second principle, combination heredity, asserts that a combination is more likely to be failure-inducing when some of its component inputs are more likely to be failure-inducing. From our product-form prior (2), note that the combination root cause probability $p_{(\mathbf{i},\mathbf{j})_{K}}$ cannot be large unless some of its component root cause probabilities in $\{p_{(i_{k},j_{k})}\}_{k=1}^{K}$ are also large. This thus captures the desired combination heredity effect. Similar product-form weights have been used for modeling hierarchy and heredity in the context of predictive modeling (Tang et al.,, 2023) and data reduction (Mak and Joseph,, 2017).
@@
With the product-form prior (2), we require only the specification of the single-input root cause probabilities $\{p_{(i,j)}\}_{i,j}$, which can be used to integrate further domain knowledge from test engineers. For most software systems at the testing stage, it may be reasonable to specify a small (i.e., near-zero) value for $p_{(i,j)}$, as this reflects the prior belief that failure-inducing root causes should occur sporadically. Oftentimes, however, an engineer has additional domain knowledge that permits a more informed prior specification, e.g., from historical data or expert judgment. For example, the engineer may know that certain factors have been recently added to the system, and thus may be more suspicious of such factors. This heightened suspicion can be captured via a larger specification of its $p_{(i,j)}$ compared to other factors. We shall see how such domain knowledge can accelerate fault localization in later case studies.
@@
From a Bayesian perspective, the product-form prior (2) provides a way for propagating elicited domain knowledge over the many root cause probabilities in $\mathbf{p}$. For example, suppose an engineer has heightened suspicions on factor $i$, and accordingly specifies a higher value for the single-factor root cause probabilities $\{p_{(i,j)}\}_{j}$. By (2), this induces larger prior root cause probabilities $p_{(\mathbf{i},\mathbf{j})_{K}}$ for any combination $(\mathbf{i},\mathbf{j})_{K}$ involving factor $i$, thus “pooling” this information over such combinations. This “information pooling”, guided by the embedded principles of combination hierarchy and heredity, can facilitate the disentangling of the large number of potential root causes from limited test runs. Recent work on related notions of information pooling have shown promise in high-dimensional inference problems, e.g., matrix completion (Yuchi et al.,, 2023) and multi-armed bandits (Mak et al.,, 2022); we show that this is also similarly important for timely fault localization. Of course, as with any Bayesian modeling framework, the domain and/or structural knowledge captured within the BayesFLo prior is not a strict assumption; it can be overturned with sufficient evidence from data.
@@ 3.2 Posterior Root Cause Probabilities
Figure 3 (left) visualizes this partition of $\mathcal{C}$ from observed test runs for a simple example.
Figure 3: [Left] Visualizing the use of passed and failed test cases for partitioning the set of considered combinations $\mathcal{C}$ into $\mathcal{C}_{\rm TP}$, $\mathcal{C}_{\rm TF}$ and $\mathcal{C}_{\rm UT}$. [Right] Workflow for the proposed BayesFLo fault localization approach.
@@
$\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})=0,\quad(\mathbf{i},\mathbf{j})\in\mathcal{C}_{\rm TP}.$ (5)
This is akin to Ghandehari et al., (2018), which removes TP combinations from consideration for root causes. Furthermore, for UT combinations, we have:
@@
since the observed test set $\mathcal{D}$ does not provide any information on an untested combination $(\mathbf{i},\mathbf{j})$. As such, its root cause probability given $\mathcal{D}$ simply reduces to its prior probability given in (2). The challenge thus lies in computing posterior probabilities on the remaining class of TF combinations. We present next an approach for computing such probabilities, leveraging tools from integer programming and graph representations. Figure 3 (right) summarizes the proposed algorithmic workflow; we elaborate on this in the following section.
4 Computation of Root Cause Probabilities
@@
where $\mathbb{P}(\mathbf{Z}=\mathbf{z})$ follows from Equation (2), and $\mathbb{P}(\mathcal{D}|\mathbf{Z}=\mathbf{z})$ can be deduced from the reasoning in Section 3.2. The limitation of such an approach is clear. For each $(\mathbf{i},\mathbf{j})\in\mathcal{C}_{\rm TF}$, we need to compute the sum of $2^{|\mathcal{C}|-1}$ terms in the numerator and the sum of $2^{|\mathcal{C}|}$ terms in the denominator. Hence, even for small systems with $|\mathcal{C}|$ small, this brute-force approach can be infeasible. This sheer dimensionality of potential root cause scenarios is the key bottleneck for tractable computation of probabilities for Bayesian fault localization.
To address this, we employ an alternate formulation, which permits considerable speed-ups for computing probabilities. We first outline this reformulation, then show how this facilitates efficient computation via a connection to the related problem of minimal set covering.

### v3

@@ 2.2 State-of-the-Art and Its Limitations
Section 1 provides a literature review of existing fault localization methods, which we do not repeat here for brevity. We review next two state-of-the-art techniques, the BEN approach in Ghandehari et al. (2020) and the weighting approach in Lekivetz and Morgan (2018), which will serve as our benchmarks later.
The BEN approach in Ghandehari et al. (2020) is a two-phase method, where Phase 1 identifies and ranks potential root causes and Phase 2 locates faulty statements. Phase 1 is most relevant for our fault localization goal, as Phase 2 considers the downstream problem of debugging the software system. Phase 1 of BEN employs the so-called component suspiciousness metric, defined for a component $(i,j)$ (i.e., level $j$ of factor $i$) as $\rho(i,j)=\frac{1}{3}[u(i,j)+v(i,j)+w(i,j)]$. Here, $u(i,j)$ denotes the ratio of failed test cases with component $(i,j)$ over the total number of failed test cases, $v(i,j)$ denotes the ratio of failed test cases with $(i,j)$ over the total number of test cases with $(i,j)$, and $w(i,j)$ denotes the ratio of suspicious combinations with $(i,j)$ over the total number of suspicious combinations. The suspiciousness metric of an input combination is defined as the average component suspiciousness of its components. This metric is then used (along with a measure of environmental suspiciousness) to rank the suspiciousness of potential root causes from test outcomes.
One drawback with BEN is that it cannot incorporate domain knowledge from test engineers. To address this, Lekivetz and Morgan (2018) adopts an alternate weighting approach. Here, a weight is first assigned to each factor level: the baseline weight is 1, and a higher weight is assigned if one believes (e.g., from expert judgment or system characteristics) that such a level is more likely to cause a failure. The weight of an input combination is taken as the product of the weights for its components. Suspicious combinations are then ranked based on their “normalized” weights. In the simple case where each failed run involves a separate set of potential root causes, the normalized weight of a suspicious combination is computed as its weight divided by the sum of weights for all potential root causes for its corresponding failed run. When there are overlaps in potential root causes between different failed runs, a complete enumeration of potential root cause scenarios is needed to compute the normalized weights. This enumeration is possible via brute force with a small number of failed test runs. However, as the number of failed runs grows, such an enumeration quickly becomes computationally infeasible; more on this in Section 5.2.
Note that, for this weighting approach, the weight of a higher-order combination (e.g., factors A, B, C at level 1) is by construction larger than that for a lower-order combination (e.g., factors A, B at level 1). Thus, applying this weighting approach for combinations of different orders may lead to unintended heightened suspiciousness of higher-order combinations, which is undesired. To avoid this, Lekivetz and Morgan (2018) recommends that this approach be used to iteratively analyze combinations of the same order, beginning with lower-order combinations (which have greater suspiciousness a priori by the principle of combination hierarchy; see Lekivetz and Morgan, 2021). In other words, combinations of order 2 should be analyzed and debugged first, then combinations of order 3, and so on. The same iterative analysis procedure over combination order is employed for the BEN approach in Ghandehari et al. (2020) for a similar reason.
We now investigate these methods for the motivating TCAS application. Here, the true root cause – a fully-documented fault in Griesmayer et al. (2007) – is a single 3-factor combination involving factors 12, 8 and 9; this fault is further discussed in Section 5.1. The BEN approach in Ghandehari et al. (2020) is implemented following the algorithm in their paper. For the weighting approach in Lekivetz and Morgan (2018), we incorporate domain knowledge on the nature of the underlying fault via higher weights (i.e., heightened suspiciousness) on certain factors; details are provided later in Section 5.1. Since there are only two failed test runs here, its normalized weights can be computed via a brute-force enumeration of potential root cause scenarios. Table 3 shows the ranked list of 2-factor and 3-factor combinations (2-FCs and 3-FCs) identified by these two approaches, with the true 3-factor root cause bolded.
This table reveals two notable limitations of these existing methods. First, since both methods employ an iterative analysis approach with respect to combination order, they require the investigation of all eight 2-FCs before finding the true 3-factor root cause. Given the expensive nature of debugging each combination (which can take thousands of dollars), this considerably increases the cost of identifying the true root cause. A key goal is thus to identify the true root cause with as few such follow-up investigations as possible (Landsberg and Barr, 2018). A Bayesian approach can achieve this via a probabilistic modeling framework over all combination orders, which permits a fair comparison of suspicious combinations across orders via its posterior probabilities. Second, while both approaches employ different suspiciousness metrics for ranking combinations, such metrics are not posterior probabilities and do not provide a principled quantification of risk for each combination. This can hinder actionable decisions: here, it is not clear from existing analyses whether one should, e.g., investigate all 3-FCs, or whether one can confidently ignore the lower-ranked combinations. A Bayesian approach can address this limitation, by providing posterior probabilities on suspicious combinations (along with a user-specified significance level) that can help plan the debugging effort needed for confident fault localization.
Combination Type Count Normalized Weight
2-FC 1 1.00
2-FC 2 0.22
2-FC 5 0.11
3-FC 1 0.65
3-FC 140 << 0.02
(a) Weighting approach.
Combination Type Count Rank
2-FC 2 1
2-FC 6 3-8
3-FC 1 1
3-FC 140 2-141
(b) BEN (Phase 1).
Table 3: Suspicious 2-FCs and 3-FCs ranked using the weighting approach in Lekivetz and Morgan (2018) and the BEN approach (Phase 1) in Ghandehari et al. (2020) for our motivating TCAS case study. Combinations at the top are deemed more suspicious and are investigated first. The true root cause is bolded.
@@ 3 The BayesFLo Model
We propose a new Bayesian Fault Localization (BayesFLo) framework, which provides a principled probabilistic approach for fault localization via conditioning on test set outcomes. We first present this Bayesian model, then show how it permits the integration of domain and structural knowledge within its prior specification.
Here, it is worth mentioning that, while there may be different “types” of failures (e.g., incorrect or missing outputs), the system has typically undergone a reasonable level of preliminary testing (e.g., unit testing, which eliminates root causes in single units) at time of combinatorial testing. Because of this, there remains only a very limited number of faults for diagnosis (Lekivetz and Morgan, 2021). As such, we presume in the following only a single failure type at time of BayesFLo analysis. We will briefly discuss an approach for the rare scenario of multiple failure types later in the Conclusion.
@@ 3.1 Prior Specification
We first introduce some notation. Consider a software system (or more broadly, a complex engineering system) with $I\geq 1$ input factors, where a factor $i$ can take on $J_{i}\geq 2$ different levels. A $K$-factor combination (with $K\leq I$) is denoted as $(\mathbf{i},\mathbf{j})_{K}$, where $\mathbf{i}=(i_{1},\cdots,i_{K}),i_{1}<\cdots<i_{K}$ is an ordered $K$-vector containing all inputs for this combination, and $\mathbf{j}=(j_{1},\cdots,j_{K}),j_{k}\in\{1,\cdots,J_{i_{k}}\}$ is a $K$-vector indicating the levels of each corresponding input. For example, the 2-factor combination of the first factor at level 1 and the second factor at level 2 can be denoted as $(\mathbf{i},\mathbf{j})_{2}$, where $\mathbf{i}=(1,2)$ and $\mathbf{j}=(1,2)$. In the case of $K=1$, i.e., a single input $i$ at level $j$, this may be simplified to $(i,j)$.
Now let $\mathcal{C}_{K}$ denote the set of $K$-factor combinations $(\mathbf{i},\mathbf{j})_{K}$ as described above, and let $\mathcal{C}=\cup_{K=1}^{I}\mathcal{C}_{K}$ denote the set of combinations over all orders $K=1,\cdots,I$. Further let $Z_{(\mathbf{i},\mathbf{j})_{K}}\in\{0,1\}$ be an indicator variable for whether the combination $(\mathbf{i},\mathbf{j})_{K}$ is truly a root cause. As this is unknown prior to running test cases, we model each $Z_{(\mathbf{i},\mathbf{j})_{K}}$ a priori as an independent Bernoulli random variable of the form:
@@
It is worth noting the sheer number of input combinations in $\mathcal{C}$ that needs to be considered as potential root causes. Assuming each factor has an equal number of levels $J=J_{1}=\cdots=J_{I}$, one can show that $\mathcal{C}_{K}$ contains $\binom{I}{K}J^{K}$ distinct combinations of order $K$, thus the total number of considered input combinations is $|\mathcal{C}|=$ $\sum_{K=1}^{I}\binom{I}{K}J^{K}$. Even with a moderate number of inputs, say $I=10$, with each having $J=2$ levels, this amounts to $|\mathcal{C}|=59,048$ combinations. As we shall see later in Section 4, the size of $\mathcal{C}$ forms the key bottleneck for Bayesian inference, as the computation of posterior probabilities can require $\mathcal{O}(2^{|C|})$ work; this can thus be infeasible even for small software systems.
@@
In words, the combination root cause probability $p_{(\mathbf{i},\mathbf{j})_{K}}$ is modeled as the product of the root cause probabilities for its component inputs. The key advantage of this form is that it captures structural knowledge on root causes, by embedding the principles of combination hierarchy and heredity (Lekivetz and Morgan, 2021). These principles reflect the structured nature of typical software root causes, and can be seen as extensions of the well-known principles of effect hierarchy and heredity (Wu and Hamada, 2009), which are widely used for analysis of factorial experiments. The first principle, combination hierarchy, asserts that combinations involving fewer inputs are more likely to be failure-inducing than those involving more inputs. Empirical evidence suggests this principle holds for software systems across various domains (Kuhn et al., 2004). To see how our prior (2) captures combination hierarchy, note that by its product form construction, the combination probability $p_{(\mathbf{i},\mathbf{j})_{K}}$ is always less than the probability of any component input $p_{(i_{k},j_{k})}$. Thus, this prior assigns increasingly smaller root cause probabilities on combinations with a higher combination order $K$, thus capturing the desired hierarchy structure. The second principle, combination heredity, asserts that a combination is more likely to be failure-inducing when some of its component inputs are more likely to be failure-inducing. From our product-form prior (2), note that the combination root cause probability $p_{(\mathbf{i},\mathbf{j})_{K}}$ cannot be large unless some of its component root cause probabilities in $\{p_{(i_{k},j_{k})}\}_{k=1}^{K}$ are also large. This thus captures the desired combination heredity effect. Similar product-form weights have been used for capturing hierarchy and heredity for predictive modeling (Tang et al., 2023) and data reduction (Mak and Joseph, 2017).
@@
With the product-form prior (2), we require only the specification of the single-input root cause probabilities $\{p_{(i,j)}\}_{i,j}$, which can be used to integrate further domain knowledge from test engineers. For most software systems at the testing stage, it may be reasonable to specify a small (i.e., near-zero) value for $p_{(i,j)}$, as this reflects the prior belief that failure-inducing root causes should occur sporadically. Oftentimes, however, an engineer has additional domain knowledge that permits a more informed prior specification, e.g., from historical data or expert judgment. For example, the engineer may know that certain factors have been recently added to the system, and thus may be more suspicious of such factors. This heightened suspicion can be captured via a larger specification of its $p_{(i,j)}$ compared to other factors. We shall see how such domain knowledge can improve fault localization in later case studies.
@@
From a Bayesian perspective, the product-form prior (2) provides a way for propagating the elicited domain knowledge over the many root cause probabilities in $\mathbf{p}$. For example, suppose an engineer has heightened suspicions on factor $i$, and accordingly specifies a higher value for the single-factor prior probabilities $\{p_{(i,j)}\}_{j}$. By (2), this induces larger prior root cause probabilities $p_{(\mathbf{i},\mathbf{j})_{K}}$ for any combination $(\mathbf{i},\mathbf{j})_{K}$ involving factor $i$, thus “pooling” this information over such combinations. This “information pooling”, guided by the embedded principles of combination hierarchy and heredity, can facilitate the disentangling of the large number of potential root causes from limited test runs. Recent work on related notions of information pooling have shown promise in high-dimensional inference problems, e.g., matrix completion (Yuchi et al., 2023) and multi-armed bandits (Mak et al., 2022); we show that this is also similarly important for efficient fault localization. Of course, as with any Bayesian modeling framework, the domain and/or structural knowledge captured within the BayesFLo prior is not a strict assumption; it can be overturned with sufficient evidence from data.
@@ 3.2 Posterior Root Cause Probabilities
Figure 3 visualizes this partition of $\mathcal{C}$ from observed test runs for a simple example.
Figure 3: Visualizing the use of passed and failed test cases for partitioning the set of considered combinations $\mathcal{C}$ into $\mathcal{C}_{\rm TP}$, $\mathcal{C}_{\rm TF}$ and $\mathcal{C}_{\rm UT}$.
@@
$\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})=0,\;(\mathbf{i},\mathbf{j})\in\mathcal{C}_{\rm TP}.$ (5)
This is akin to Ghandehari et al. (2020), which removes TP combinations from consideration for root causes. Furthermore, for UT combinations, we have:
@@
since the observed test set $\mathcal{D}$ does not provide any information on an untested combination $(\mathbf{i},\mathbf{j})$. As such, its root cause probability given $\mathcal{D}$ simply reduces to its prior probability given in (2). The challenge thus lies in computing posterior probabilities on the remaining class of TF combinations. We present next an approach for computing such probabilities, leveraging tools from integer programming and graph representations.
4 Computation of Posterior Root Cause Probabilities
@@
where $\mathbb{P}(\mathbf{Z}=\mathbf{z})$ follows from Equation (2), and $\mathbb{P}(\mathcal{D}|\mathbf{Z}=\mathbf{z})$ can be deduced from the reasoning in Section 3.2. The limitation of such an approach is clear. For each $(\mathbf{i},\mathbf{j})\in\mathcal{C}_{\rm TF}$, we need to compute the sum of $2^{|\mathcal{C}|-1}$ terms in the numerator and the sum of $2^{|\mathcal{C}|}$ terms in the denominator. Hence, even for small systems with $|\mathcal{C}|$ small, this brute-force approach can be infeasible. This sheer dimensionality of potential root cause scenarios is a key bottleneck for tractable computation of posterior probabilities for Bayesian fault localization.
To address this, we employ an alternate reformulation, which permits considerable speed-ups for computing probabilities. We first outline this reformulation, then show how this facilitates efficient computation via a connection to the related problem of minimal set covering.